# Autonomous Robotic Bone Micro-Milling System with Automatic Calibration and 3D Surface Fitting

### v2

@@ V-D3 Result and discussion
V-E Experiment 3: Validation of the autonomous robotic bone micro-milling system
Figure 8: The automation flow of an experimental case. (a) The result of Automatic Calibration (Section III-A). (b) The result of 3D Surface Fitting (Section III-B). (c) The input and output of the Completion Level Recognition (Section IV-A). (d) The result after the milling process was completed and after the bone flap was manually removed.
Following the end-to-end automation workflow shown in Fig. 2, with its hardware setup shown in Fig. 1 and described in Section II, autonomous cranial window creation experiments were conducted on euthanized mice to validate the feasibility of the proposed method.
V-E1 Animal model
@@ V-E2 Metrics
Milling success and time are used to evaluate the performance. It should be noted that milling success is defined as the bone flap (see Fig. 8-(d)) within the milling trajectory being removable manually with tweezers after the milling is completed, without damaging the membrane beneath the skull [8].
V-E3 Result and discussion
7 experiments were conducted, and 6 were successful, resulting in a success ratio of 85.7 %. The average required time for successful milling was $2.1\pm 1.8$ minutes. Compared with the 80 % of success ratio and the $16.8\pm 2.5$ minutes of milling time using the previous system [8] evaluated on eggshell model, the feasibility and efficiency of the improved autonomous robotic bone micro-milling scheme were initially proved on euthanized mice. An example of the entire flow of a successful milling case is shown in Fig. 8. The improvement in success ratio and the reduction in milling time aligned with our expectations, as the introduction of MSCS enhanced the system’s perceptual capability for uneven surfaces, avoiding the long milling time and potential failures caused by the complete reliance on real-time feedback in the previous system.
@@ V-E3 Result and discussion
Furthermore, compared with the results of teleoperated robotic cranial window creation mentioned in Section V-C ($13.7\pm 6.6$ minutes, 3 skilled and 4 novice human operators, 10 euthanized mice) and manual cranial window creation [6] ($8.0\pm 2.1$ minutes, 2 skilled human operators, 6 anesthetized live mice), the system is demonstrated with excellent speed, repeatability, and consistency.
The error in Completion Level Recognition is considered the main cause of the failure case, indicating that the recognition accuracy of the CNN relying solely on RGB images limits the upper bound of the success ratio. Future work will require improving the recognition accuracy of the CNN or adding additional sensors to assist with image processing.
VI Discussion and Conclusions
@@
In this paper, we proposed a pre-measurement pipeline of automatic calibration and 3D surface fitting, and integrate it into an autonomous robotic bone micro-milling system, enabling an end-to-end autonomous cranial window creation workflow capable of rapidly, in real-time, and accurately perceiving and adapting to the uneven surface and non-uniform thickness of bone tissue without human assistance. In the experiments, we validated the system’s excellent performance and efficiency by successfully performing autonomous cranial window creation on euthanized mice, achieving faster, more accurate and more stable results compared to skilled human operators.
Future works include further improving the performance of the system and applying the system on anesthetized live mice. In addition, considering the similar challenges of target surface and thickness uncertainties in many surgeries, such as endoscopic endonasal surgery and neural surgery, we also hope to apply our autonomous robotic bone micro-milling system to other medical fields.
@@ Acknowledgment
We would like to express our special gratitude to Dr. Saúl Alexis Heredia Pérez, Dr. Murilo Marques Marinho and the surgeons in Institute of Science Tokyo for their precious advice and kind help.

### v3

@@ IV-C3 Result and discussion
IV-D Experiment 3: Validation of the autonomous robotic bone micro-milling system
Figure 8: The automation flow of an experimental case. (a) The result of Automatic Calibration (Section II-A). (b) The result of 3D Surface Fitting (Section II-B). (c) The input and output of the Completion Level Recognition (Section III-A). (d) The result after the milling process was completed and after the bone flap was manually removed.
Following the end-to-end automation workflow shown in Fig. 2, with its hardware setup shown in Fig. 1, autonomous cranial window creation experiments were conducted on euthanized mice to validate the feasibility of the proposed method.
IV-D1 Animal model
@@ IV-D2 Metrics
Milling success and time are used to evaluate the performance. Success is defined as the bone flap (see Fig. 8-(d)) within the milling trajectory being removable manually with tweezers after the milling is completed, without damaging the membrane beneath the skull [21].
IV-D3 Result and discussion
7 experiments were conducted, and 6 were successful, resulting in a success rate of 85.7 %. The average required time for successful milling was $2.1\pm 1.8$ minutes. Compared with the 80 % of success rate and the $16.8\pm 2.5$ minutes of milling time using the previous system [21] evaluated on the eggshell model, the feasibility and efficiency of the improved autonomous robotic bone micro-milling scheme were initially proved on euthanized mice. The overall improvement in milling performance highlights the importance of the proposed online pre-measurement pipeline, which effectively eliminates idle time caused by the lack of pre-measured anatomical information of the skull. Crucially, this substantial performance gain was achieved in the more challenging context of the non-uniform, biological mouse skull, in contrast to the simpler, uniform eggshell model used in the prior study. Fig. 8 shows a successful milling case. As expected, the MSCS improved surface perception, increasing the success rate and reducing milling time compared to the previous system that relied solely on real-time feedback.
@@ IV-D3 Result and discussion
Furthermore, compared with the results of teleoperated robotic cranial window creation mentioned in Section IV-B ($13.7\pm 6.6$ minutes, 3 skilled and 4 novice human operators, 10 euthanized mice) and manual cranial window creation [13] ($8.0\pm 2.1$ minutes, 2 skilled human operators, 6 anesthetized live mice), the system is demonstrated with excellent efficiency, repeatability, and consistency.
In the 7 experiments, the only failure case was primarily caused by an error in the Completion Level Recognition. In this case, the completion level of the area that had already been completely penetrated, which should have been 1, was incorrectly recognized as being below 1, causing the drill tip to continue descending. Consequently, the underlying membrane was damaged and the automatic milling process was stopped manually. We consider that the accuracy of the current recognition CNN, relying solely on RGB images, limits the upper bound of the system’s success rate. Future work will focus on improving the recognition accuracy of the CNN or incorporating additional signals to assist image processing.
It is noteworthy that a potential failure mode in previous systems did not occur in our experiments, where prolonged milling duration caused the drill to repeatedly contact already penetrated membrane areas, eventually leading to rupture. By shortening the overall milling duration, we effectively reduced the drill’s dwell time on these vulnerable areas, thereby lowering the risk of membrane damage.
V Discussion and Conclusions
@@
In this paper, we propose an online pre-measurement pipeline of automatic calibration and 3D surface fitting, and integrate it into an autonomous robotic bone micro-milling system, enabling an end-to-end autonomous cranial window creation workflow capable of rapidly, in real-time, and accurately perceiving and adapting to the uneven surface and non-uniform thickness of bone tissue. Crucially, this work goes beyond the framework of the previous system by eliminating the need for manual assistance and addressing the main limitations, resulting in a system that is not only truly autonomous but also achieves comprehensive improvements in performance and efficiency. In the experiments, we validated the system’s excellent performance and efficiency by successfully performing autonomous cranial window creation on euthanized mice, achieving faster, more accurate and more stable results compared to skilled human operators.
Future work includes further improving the system performance and applying it to anesthetized live mice. Beyond the mouse cranial window creation, the proposed framework holds strong potential for generalization to other microsurgical bone milling tasks requiring high precision and tissue preservation. The integration of pre-measurement and real-time feedback enables adaptive control over heterogeneous biological structures, suggesting a general robotic strategy for microscale bone processing and possible extensions to broader medical applications such as endoscopic or neurosurgical procedures.
@@ Acknowledgment
We express our special gratitude to Dr. Saúl Alexis Heredia Pérez, Dr. Murilo Marques Marinho, and the surgeons at the Institute of Science Tokyo for their advice and help.